# Inference of time delay in stochastic systems

### v1

@@ III.1 Definitions
The power spectral density (PSD) [54, 55] is a well established quantity for the analysis of noisy (measurement noise) or fully stochastic (e.g., Brownian) systems. It is frequently used in experimental and theoretical studies [39, 40, 41, 42, 56, 57, 58, 59] as well as a wide range of (engineering) applications (e.g., signal analysis [55, 60]). Its main advantages are accessibility (i.e., it can be directly applied to measurement data and computed using fast algorithms [61] such as the fast Fourier transform (FFT) [62]) and robustness with respect to noise.
In the context of linear time-delayed feedback, the PSD has previously been computed using correlation functions [52] and as a result derived from response theory [63, 64] (see also Appendix B.1). We here focus on a different definition (often used in signal analysis [55]), it reads [65]
@@
where $\mathcal{T}$ is the trajectory length, $\tilde{Z}$ is the Fourier transform of a stationary time series $Z(t)$ (i.e., with mean and variance approaching a constant value) and $\omega$ is the frequency. Here, we used the following definitions for the Fourier transform:
[…]
For a stochastic time series $Z(t)$, the power spectral density (5) is a fluctuating quantity. To ensure that the data are suitably smooth for the purpose of inference of the delay time, we consider the noise-averaged power spectral density (from this point we use the acronym PSD as a short hand notation for this quantity), it is defined as
@@
[…]
@@ III.2 Linear systems
Figure 1: Power spectral density (PSD) as a function of the frequency $\omega$ for a Brownian particle subject to linear time-delayed feedback (2) at $\tau=0.15\tau_{B}$. Solid/dashed lines: Analytical result (11) (panel a) and the PSD multiplied by $\omega^{2}$ (panel b), for different values of $k_{a,b}$. Black solid curve represents a harmonically confined Brownian particle without feedback [i.e., $k_{b}=0$ in Eq. (11). Dash-dotted grey vertical lines: Approximate inflection points at $\omega=2\pi n/\tau$ with $n\in\mathbb{N}$. Dotted black vertical lines: Approximating maximum for $k_{a,b}>0$ at $\omega=3\pi/(2\tau)+2\pi n/\tau$ for $n>1$, where $n\in\mathbb{N}$.
We first consider the limit $k_{b}=0$ in Eq. (11), i.e., no feedback. In this case, and if $k_{a}>0$, the expression takes the shape of the PSD of a Brownian particle in a harmonic trap of stiffness $k_{a}$, i.e., $\langle S(\omega)\rangle_{\mathrm{lin}}=\frac{2\gamma k_{B}T}{\gamma^{2}\omega^{2}+k_{a}^{2}}$, which decreases as $1/\omega^{2}$ in the large-frequency limit (see the non-oscillatory black solid line in Fig. 1). (We note here that for $k_{a}<0$, the particle is subject to an instantaneous repulsive harmonic potential, which yields instability.)
When introducing time delay, i.e., $\tau>0$, the $\tau$ dependent oscillatory terms ($\sin$ and $\cos$) in the denominator of Eq. (11) come into play. As a consequence, the PSD and even more clearly the function $\omega^{2}$ [that is, Eq. (11) multiplied by $\omega^{2}$] show oscillatory behavior, in contrast to the harmonically trapped particle (see Fig. 1). Oscillations induced by time delay are not unexpected. In deterministic systems, this can be an oscillating position while for stochastic systems one would expect oscillations in the mean square displacement [49] or other correlation functions [12, 63, 64]. Furthermore, for similar linear feedback force systems, we know that at least in stationary states, oscillations can also occur in the PSD.
@@
Importantly, the specific choice of the parameters $k_{a,b}$, only influences the small frequency behavior (see Appendix B.2 for an analytical consideration of this limit) but does not affect the oscillatory part of the PSD on which we concentrate from now on.
@@
Figure 2: Power spectral density (PSD) as a function of the frequency $\omega$ for a Brownian particle subject to linear time-delayed feedback (2) for $k_{b}=0.5k_{a}>0$. Solid lines: Analytical results (11) (panel a) and the PSD multiplied by $\omega^{2}$ (panel b) for different values of the delay time $\tau$. Dash-dotted vertical lines: Approximate inflection points at $\omega=2\pi n/\tau$ with $n\in\mathbb{N}$. Dotted vertical lines: Approximating maximum at $\omega=3\pi/(2\tau)+2\pi n/\tau$ for $n>1$, where $n\in\mathbb{N}$.
@@
To illustrate the $\tau$-dependence of the oscillations, we present the PSD and $\omega^{2}\mathrm{PSD}$ at different values of the delay time (color coded) in Fig. 2 for a system where $k_{a}>k_{b}>0$. While the term $2k_{a}k_{b}\cos(\omega\tau)$ in Eq. (11) has a maximum magnitude of $|2k_{a}k_{b}|$, the magnitude of $2\gamma k_{b}\omega\sin(\omega\tau)$ grows with $\omega$. Therefore, as $\omega$ grows, specifically when $\omega>k_{a}/\gamma$, the term $2\gamma k_{b}\omega\sin(\omega\tau)$ [in Eq. (11)] becomes dominant. Indeed, we observe from Fig. 2 that the value of the delay time determines both, the location of the maxima and minima and the oscillation period. The first minimum/maximum after the sine term becomes dominant is the first relevant maximum/minimum for attractive/repulsive feedback. It is located at approximately $\omega=3\pi/(2\tau)$. For $\omega\gg k_{a}/\gamma$, the oscillation period [due to the periodicity of the $\sin$ and $\cos$ terms in Eq. (11)] is given by $\Delta\omega=2\pi/\tau$ in frequency space (dotted vertical lines of the same color indicating the location of the maxima and inflection points, respectively).
@@
For longer delay times, we observe from Fig. 2 a shift of the delay-time dependent locations of the minima and maxima to smaller frequencies and a shorter oscillation period, consistent with the prediction. The signature of time delay, i.e., the specific oscillatory behavior and delay-time dependent oscillation period $\Delta\omega=2\pi/\tau$ can, in principle, be used to infer time delay without exact knowledge about the other parameters in the system or fitting simulation data with the analytical result.
@@
To extract the location of the maxima and minima in the frequency space, it is instructive to consider the derivative of $\omega^{2}{\rm PSD}$ with respect to $\omega$. This is depicted in Fig. 3 for the case $k_{a,b}>0$. The zero crossings correspond to the positions of the minima and maxima. For $k_{a,b}>0$ the maxima of $\omega^{2}\mathrm{PSD}$, which are roots of $\partial_{\omega}[\omega^{2}{\rm PSD}]$, are located at $\omega\approx 3\pi/(2\tau)+2\pi n/\tau$ (dotted vertical lines in Fig. 3) while the minima are offset by $\pi/\tau$. The inflection points of $\omega^{2}\mathrm{PSD}$ (maxima in Fig. 3) are located at $\omega\approx 2\pi n/\tau$, $n\in\mathbb{N}^{+}$ (indicated by dash-dotted vertical lines). We observe from Fig. 3 that from the second set of vertical lines at $\omega\tau_{B}\sim 75$, the approximation is in good agreement with the actual positions of the roots and maxima of $\partial_{\omega}[\omega^{2}{\rm PSD}]$, while for the first zero crossing and maximum there is still a significant deviation. At those frequencies, the term containing $\cos(\omega\tau)$ is not negligible.
@@
Figure 3: The derivative of PSD (11) [i.e., $\partial/\partial\omega[\omega^{2}{\rm PSD}(\omega)$] as a function of $\omega$. Here, $k_{a,b}>0$ and $k_{a}\neq k_{b}$. Dotted horizontal line: A guide for the eye to indicate the zero-crossings of the blue solid curve. Dash-dotted vertical line: Inflection points of $\omega^{2}{\rm PSD}$ at $\omega=2\pi n/\tau$, for $n\in\mathbb{N}$. Dotted vertical line: maxima of $\omega^{2}{\rm PSD}$ at $\omega=3\pi/(2\tau)+2\pi n/\tau$.
@@ IV Power spectral density approach – numerical results
We now discuss numerical results for the PSD for various types of SDDEs. We first benchmark the numerics by considering strictly linear systems. We then proceed by considering the specific nonlinear systems introduced in Sec. II.2, where linear approximations recover harmonic attractive/repulsive feedback systems. This allows us to obtain first insights into the impact of nonlinearity. Finally, we move on to more general nonlinear systems to see how the oscillations can be used to infer the delay time without explicit knowledge of the functional form of the underlying forces.
@@ IV.1 Preliminary remarks
For the computation of the PSD it is crucial that the underlying time series is stationary (see Appendix A.2 for detailed discussion on this topic), which means that for the mean and variance of $Z(t)$ in Eq. (8) have to approach constants as a function of time. Already a simple passive Brownian particle does not fulfill this requirement, since its positional variance grows in time. Similarly, the marginally stable linear feedback system effectively exhibits diffusive behavior in the long time limit. To overcome this problem, we apply the differencing method (see Appendix A.2), i.e., we compute the PSD of the positional difference time series $\{\bar{x}(t)\}\equiv\{x(t+dt)-\{x(t)\}$ for small $dt$, instead of the position itself. As a rough analogy, this method can be understood as ‘taking the discretized derivative’ of the positional data with respect to time. After applying the Fourier transform and dividing by $dt$ as well as by the frequency (recall that derivatives become algebraic factors under the Fourier transform), this leaves us with the correct Fourier transform of the position, which can then be plugged into Eq. (8). Not applying this technique would lead to a significant deviation between numerical results and analytical prediction, which become particularly severe for the (non-stationary) repulsive feedback case, as shown in Appendix A.2, Fig. A1. For consistency, we apply the same procedure to all our data, even for systems that are stationary in the long time limit.
@@
Figure 5: Comparison of numerically simulated (colored points, obtained from averaging over 400 noise realizations) power spectral density (PSD) of nonlinear systems with the linear approximation (solid/dashed curves). Panel (a): $\tanh$ feedback force for various $\tau$; solid curves: Linear approximation with $k_{a}=k_{b}=10k_{\mathrm{B}}T/\sigma$. Panel (b): $\tanh$ feedback force (3) for different $A$; solid curves: Linear approximation with corresponding $k_{a}=k_{b}>0$. Panel (c): Gaussian feedback force (4); solid (dashed) curves: Analytical result for the linear approximation with repulsive harmonic feedback ($k_{a}=k_{b}<0$) and modified attractive harmonic feedback for the constant velocity state [see Appendix B.4.2, Eq. (29) for more details]. Panel (a): $B=1/\sigma$. Panels (b) and (c): $\tau=0.15\tau_{\mathrm{B}}$. Vertical grey dotted lines: Approximate locations of maxima/minima according to the analytical result for the linear system in panels (b) and (c).
@@ IV.2 Linear systems
Figure 4 presents numerical results based on Eq. (1) with a feedback force as in Eq. (2) in the absence of additional instantaneous forces. As before, we distinguish attractive feedback (Fig. 4a) and repulsive feedback (Fig. 4b). In the regime $\omega\tau_{B}\gtrsim 1$, we find almost perfect agreement with the analytical prediction for all cases considered. Only at small frequencies and $k_{a}\neq k_{b}$ we observe a deviation from the analytical predictions. The deviation can be shifted to smaller frequencies by using longer trajectories. We therefore understand this behavior as an effect connected to the finite length of the trajectory we use to compute the PSD.
@@ IV.3 Nonlinear feedback systems and their linear approximations
Now we turn towards the nonlinear systems based on the hyperbolic tangent [Eq. (3)] and Gaussian feedback forces [Eq. (4)] that reduce to attractive/repulsive harmonic feedback systems in certain limits as discussed in Sec. III.2. Our goal here is to investigate, whether oscillatory behavior persists in nonlinear systems and under what conditions deviations from linear behavior are observed in the $\omega^{2}$. Numerical data for $\omega^{2}\mathrm{PSD}$ are presented in Fig. 5. Solid lines represent analytical results obtained for the different limits discussed in Sec. II.2 and Appendix B.4.2.
@@
First, we consider the hyperbolic tangent feedback force (3). Figure 5a and b present numerical data of the PSD and the corresponding analytical approximations based on the attractive harmonic feedback system [Eq. (2) with $k_{a}=k_{b}>0$], for different $\tau$ and $A$ respectively. For both cases, we observe good agreement between numerical data and the analytical prediction for larger $\omega$ (i.e., $\omega\tau_{\mathrm{B}}\gtrsim 10$ in Fig. 5), especially with respect to the oscillations. For small $\omega$, we observe noticeable deviations from the analytical linear prediction. These deviations grow, as $\tau$ is increased or $A$ is decreased. This is because we are approaching the limit where the linear approximation becomes invalid. Similar to linear systems (Fig. 2), for this nonlinear system the delay time determines the oscillation period as well as the location of the minima and maxima.
@@
Moving on to the system subject to Gaussian time-delayed feedback [Eq. (4)], we have to distinguish two limits based on the deterministic threshold, i.e., $a\tau/(\gamma b^{2})=1$. Far below this threshold (in particular $a=2k_{B}T$ and $a=3k_{B}T$), in the limit of small displacements the feedback force can be approximated as a repulsive harmonic feedback [50]. On the other hand, for feedback parameters far beyond the threshold to persistent motion ($A=55k_{B}T$ and $A=85k_{B}T$), the particle can be considered as moving at an approximately constant velocity while being subject to attractive harmonic feedback [50] (see Appendix B.4.2). From Fig. 5c, far above and below the threshold, we see agreement between the data obtained by numerical simulations of the full nonlinear problem and the expression Eq. (11) and the result based on (32) respectively. Particularly, we observe oscillations with the same characteristics as in the linear system. Only at low frequencies, i.e., before the onset of oscillatory behavior, do the simulation data deviate from the analytical approximation.
@@
Interestingly, as we approach the threshold, $a\tau/(\gamma b^{2})=1$, from above or below ($a=6k_{B}T$ and $a=25k_{B}T$) we see a significant deviation from the analytical prediction at small $\omega$, indicating that the simple linear approximation is no longer valid but at best a crossover behavior becomes relevant. However, the locations of the minima remain unaffected, from which we interpret that the signature of time delay remains even outside the linear regime.

### v2

@@ III.1 Definitions
The power spectral density (PSD) [66, 67] is a well established quantity for the analysis of noisy (measurement noise) or fully stochastic (e.g., Brownian) systems. It is frequently used in experimental and theoretical studies [51, 52, 53, 54, 68, 69, 70, 71, 72] as well as a wide range of (engineering) applications (e.g., signal analysis [67, 73]). Its main advantages are accessibility (i.e., it can be directly applied to measurement data and computed using fast algorithms [74] such as the fast Fourier transform (FFT) [75]) and robustness with respect to noise.
In the context of linear time-delayed feedback, the PSD has previously been defined and computed based on correlation functions [64] and in the context of response theory [76, 77] (see also Appendix A.2). We here focus on a different definition (often used in signal analysis [67]). It reads
@@
where $\mathcal{T}$ is the trajectory length, $\tilde{Z}$ is the Fourier transform of a stationary time series $Z(t)$ (i.e., with mean and variance approaching a constant value) and $\omega$ is the frequency. Here, we used the following definition for the Fourier transform:
$\tilde{f}(\omega)$ $=\int_{-\infty}^{+\infty}~dt~e^{-i\omega t}~f(t)\ .$ (6)
For a stochastic time series $Z(t)$, the power spectral density (5) is a fluctuating quantity. To ensure that the data are suitably smooth for the purpose of inference of the delay time, we consider the noise-averaged power spectral density (from this point we use the acronym PSD as a short hand notation for this quantity). It is defined as
@@
where $\langle\ldots\rangle$ indicates an ensemble average over noise realizations. To compute the PSD, we evaluate the ensemble average before taking the limit $\mathcal{T}\to\infty$. Swapping the average and the limit is possible as long as each trajectory is sufficiently long.
@@ III.2 Linear systems
Figure 1: Power spectral density (PSD) as a function of the frequency $\omega$ for a Brownian particle subject to linear time-delayed feedback (2) at $\tau=0.15\tau_{B}$. Solid/dashed lines: Analytical result (10) (panel a) and the PSD multiplied by $\omega^{2}$ (panel b), for different values of $k_{a,b}$. Black solid curve represents a harmonically confined Brownian particle without feedback [i.e., $k_{b}=0$ in Eq. (10). Dash-dotted grey vertical lines: Approximate inflection points at $\omega=2\pi n/\tau$ with $n\in\mathbb{N}$. Dotted black vertical lines: Approximate maximum for $k_{a,b}>0$ at $\omega=3\pi/(2\tau)+2\pi n/\tau$ for $n>1$, where $n\in\mathbb{N}$.
We first consider the limit $k_{b}=0$ in Eq. (10), i.e., no feedback. In this case, and if $k_{a}>0$, the expression takes the shape of the PSD of a Brownian particle in a harmonic trap of stiffness $k_{a}$. It follows that $\langle S(\omega)\rangle_{\mathrm{lin}}=\frac{2\gamma k_{B}T}{\gamma^{2}\omega^{2}+k_{a}^{2}}$, which decreases as $1/\omega^{2}$ in the large-frequency limit (see the non-oscillatory black solid curve in Fig. 1). (We note here that for $k_{a}<0$, the particle is subject to an instantaneous repulsive harmonic potential, which yields instability.)
When introducing time delay, i.e., $\tau>0$, the $\tau$-dependent oscillatory terms ($\sin$ and $\cos$) in the denominator of Eq. (10) come into play. As a consequence, the PSD and, even more clearly, the function $\omega^{2}$PSD [that is, Eq. (10) multiplied by $\omega^{2}$] show oscillatory behavior, in contrast to the harmonically trapped particle (see Fig. 1). Oscillations induced by time delay are not unexpected. In deterministic systems, this can be an oscillating position while for stochastic systems one would expect oscillations in the mean square displacement [61] or other correlation functions [12, 76, 77]. Furthermore, for similar linear feedback force systems, we know that at least in stationary states, oscillations can also occur in the PSD (see e.g. [64]). Here, we analyze the oscillations in the linear system as a reference for our subsequent investigation of the nonlinear case.
@@
Importantly, the specific choice of the parameters $k_{a,b}$, only influences the small frequency behavior (see Appendix A.3 for an analytical consideration of this limit and Appendix A.4 for the special case of a harmonic feedback system) but does not affect the oscillatory part of the PSD on which we concentrate from now on.
@@
Figure 2: Power spectral density (PSD) as a function of the frequency $\omega$ for a Brownian particle subject to linear time-delayed feedback (2) for $k_{b}=0.5k_{a}>0$. Solid lines: Analytical results (10) (panel a) and the PSD multiplied by $\omega^{2}$ (panel b) for different values of the delay time $\tau$. Dash-dotted vertical lines: Approximate inflection points at $\omega=2\pi n/\tau$ with $n\in\mathbb{N}$. Dotted vertical lines: Approximate maximum at $\omega=3\pi/(2\tau)+2\pi n/\tau$ for $n>1$, where $n\in\mathbb{N}$.
@@
To illustrate the $\tau$-dependence of the oscillations, we present the PSD and the function $\omega^{2}\mathrm{PSD}$ at different values of the delay time (color coded) in Fig. 2 for a system with $k_{a}>k_{b}>0$. While the term $2k_{a}k_{b}\cos(\omega\tau)$ in Eq. (10) has a maximum magnitude of $|2k_{a}k_{b}|$, the magnitude of $2\gamma k_{b}\omega\sin(\omega\tau)$ grows with $\omega$. Therefore, as $\omega$ grows, specifically when $\omega>k_{a}/\gamma$, the term $2\gamma k_{b}\omega\sin(\omega\tau)$ [in Eq. (10)] becomes dominant. Indeed, we observe from Fig. 2 that the value of the delay time determines both, the location of the maxima and minima and the oscillation period. The first minimum/maximum after the sine term becomes dominant is the first relevant maximum/minimum for attractive/repulsive feedback. It is located at approximately $\omega=3\pi/(2\tau)$. For $\omega\gg k_{a}/\gamma$, the oscillation period [due to the periodicity of the $\sin$ and $\cos$ terms in Eq. (10)] is given by $\Delta\omega=2\pi/\tau$ in frequency space (dotted and dash-dotted vertical lines of the same color indicating the approximate location of the maxima and inflection points, respectively). We refer to this behavior, i.e. oscillations with period $\Delta\omega=2\pi/\tau$ occurring at large $\omega$, as the signature of time delay.
@@
For longer delay times, we observe from Fig. 2 a shift of the delay-time dependent positions of the minima and maxima to smaller frequencies and a shorter oscillation period, consistent with the prediction. Thus, the specific oscillatory behavior and delay-time dependent oscillation period $\Delta\omega=2\pi/\tau$ can, in principle, be used to infer time delay without exact knowledge about the other parameters in the system or fitting simulation data with the analytical result.
@@
To extract the positions of the maxima and minima in the frequency space, it is instructive to consider the derivative of $\omega^{2}{\rm PSD}$ with respect to $\omega$. This is depicted in Fig. 3 for the case $k_{a,b}>0$. The zero crossings correspond to the positions of the minima and maxima. For $k_{a,b}>0$ the maxima of $\omega^{2}\mathrm{PSD}$, which are roots of $\partial_{\omega}[\omega^{2}{\rm PSD}]$, are located at $\omega\approx 3\pi/(2\tau)+2\pi n/\tau$ (dotted vertical lines in Fig. 3) while the minima are offset by $\pi/\tau$. The inflection points of $\omega^{2}\mathrm{PSD}$ (maxima in Fig. 3) are located at $\omega\approx 2\pi n/\tau$, $n\in\mathbb{N}^{+}$ (indicated by dash-dotted vertical lines). We observe from Fig. 3, specifically from the second set of vertical lines at $\omega\tau_{B}\sim 75$, that the approximation is in good agreement with the actual positions of the roots and maxima of $\partial_{\omega}[\omega^{2}{\rm PSD}]$, while for the first zero crossing and maximum there is still a significant deviation. At those frequencies, the term containing $\cos(\omega\tau)$ is not negligible.
@@
Figure 3: The derivative of PSD (10) [i.e., $\partial/\partial\omega[\omega^{2}{\rm PSD}(\omega)$] as a function of $\omega$. Here, $k_{a,b}>0$ and $k_{a}\neq k_{b}$. Dotted horizontal line: Guide for the eye to indicate the zero-crossings of the blue solid curve. Dash-dotted vertical line: Inflection points of $\omega^{2}{\rm PSD}$ at $\omega=2\pi n/\tau$, for $n\in\mathbb{N}$. Dotted vertical line: maxima of $\omega^{2}{\rm PSD}$ at $\omega=3\pi/(2\tau)+2\pi n/\tau$.
@@ IV Power spectral density approach – numerical results
We now discuss numerical results (based on trajectories from BD simulations, see Appendix B.1 for details) for the PSD for various types of SDDEs. We first benchmark the numerics by considering strictly linear systems. We then proceed by considering the specific nonlinear systems introduced in Sec. II.2, where linear approximations recover harmonic attractive/repulsive feedback systems. This allows us to obtain first insights into the impact of nonlinearity. Finally, we move on to more general nonlinear systems to see how the oscillations can be used to infer the delay time without explicit knowledge of the functional form of the underlying forces.
@@ IV.1 Preliminary remarks
For the computation of the PSD it is crucial that the underlying time series is stationary (see Appendix B.2 for detailed discussion on this topic), which means that for the mean and variance of $Z(t)$ in Eq. (7) have to approach constants as a function of time. Already a simple passive Brownian particle does not fulfill this requirement, since its positional variance grows in time. Similarly, the marginally stable linear feedback system effectively exhibits diffusive behavior in the long time limit. To overcome this problem, we apply the differencing method (see Appendix B.2), i.e., we compute the PSD of the positional difference time series $\{\bar{x}(t)\}\equiv\{x(t+dt)-\{x(t)\}$ for small $dt$, instead of the position itself. As a rough analogy, this method can be understood as ‘taking the discretized derivative’ of the positional data with respect to time. After applying the Fourier transform and dividing by $dt$ as well as by the frequency (recall that derivatives become algebraic factors under the Fourier transform), this leaves us with the correct Fourier transform of the position, which can then be plugged into Eq. (7). Not applying this technique would lead to a significant deviation between numerical results and analytical prediction, which become particularly severe for the (non-stationary) repulsive feedback case, as shown in Appendix B.2, Fig. B1. For consistency, we apply the same procedure to all of our data, even for systems that are stationary in the long time limit.
@@
Figure 5: Comparison of numerically simulated (colored points, obtained from averaging over 400 noise realizations) power spectral density (PSD) of nonlinear systems with the linear approximation (solid/dashed curves). Panel (a): $\tanh$ feedback force for various $\tau$; solid curves: Linear approximation with $k_{a}=k_{b}=10k_{\mathrm{B}}T/\sigma$. Panel (b): $\tanh$ feedback force (3) for different AA; solid curves: Linear approximation with corresponding $k_{a}=k_{b}>0$. Panel (c): Gaussian feedback force (4); solid (dashed) curves: Analytical result for the linear approximation with repulsive harmonic feedback ($k_{a}=k_{b}<0$) and modified attractive harmonic feedback for the constant velocity state [see Appendix A.1.2, Eq. (18) for more details]. Panel (a): $B=1/\sigma$. Panels (b) and (c): $\tau=0.15\tau_{\mathrm{B}}$. Vertical grey dotted lines: Approximate locations of maxima/minima according to the analytical result for the linear system in panels (b) and (c).
@@ IV.2 Linear systems
Figure 4 presents numerical results based on Eq. (1) with a feedback force as in Eq. (2) in the absence of additional instantaneous forces. As before, we distinguish attractive feedback (Fig. 4a) and repulsive feedback (Fig. 4b). In the regime $\omega\tau_{B}\gtrsim 1$, we find almost perfect agreement with the analytical prediction for all cases considered. Only at small frequencies and $k_{a}\neq k_{b}$ we observe a deviation from the analytical predictions. The deviation can be shifted to smaller frequencies by using longer trajectories. We therefore understand this behavior as an effect connected to the finite length of the trajectory used to compute the PSD.
@@ IV.3 Nonlinear feedback systems and their linear approximations
Now we turn towards the nonlinear systems based on the hyperbolic tangent [Eq. (3)] and Gaussian feedback forces [Eq. (4)] that reduce to attractive/repulsive harmonic feedback systems (discussed in Sec. III.2) in certain limits. Our goal here is to investigate whether the oscillatory behavior persists in nonlinear systems and under what conditions deviations from linear behavior are observed in $\omega^{2}$PSD. Numerical data for $\omega^{2}\mathrm{PSD}$ are presented in Fig. 5. Solid lines represent analytical results obtained for the different limits discussed in Sec. II.2 and Appendix A.1.2.
@@
First, we consider the hyperbolic tangent feedback force (3). Figure 5a and b present numerical data of the PSD and the corresponding analytical approximations based on the attractive harmonic feedback system [Eq. (2) with $k_{a}=k_{b}>0$], for different $\tau$ and $A$ respectively. For both cases, we observe good agreement between numerical data and the analytical prediction for larger $\omega$ (i.e., $\omega\tau_{\mathrm{B}}\gtrsim 10$ in Fig. 5), especially with respect to the oscillations. For small $\omega$, we observe noticeable deviations from the analytical linear prediction. These deviations grow as $\tau$ is increased or $A$ is decreased. This is because we are approaching the limit where the linear approximation becomes invalid. Similar to linear systems (Fig. 2), we find for this nonlinear system that the delay time determines the oscillation period, as well as the location of the minima and maxima.
@@
Moving on to the system subject to Gaussian time-delayed feedback [Eq. (4)], we have to distinguish two limits based on the deterministic threshold, i.e., $a\tau/(\gamma b^{2})=1$. Far below this threshold (in particular $a=2k_{B}T$ and $a=3k_{B}T$), in the limit of small displacements the feedback force can be approximated as a repulsive harmonic feedback [62]. On the other hand, for feedback parameters far beyond the threshold to persistent motion ($A=55k_{B}T$ and $A=85k_{B}T$), the particle can be considered as moving at an approximately constant velocity while being subject to attractive harmonic feedback [62] (see Appendix A.1.2). From Fig. 5c, far above and below the threshold, we see agreement between the data obtained by numerical simulations of the full nonlinear problem and the expression Eq. (10) and the result based on Eq. (21), respectively. Particularly, we observe oscillations with the same characteristics as in the linear system. Only at low frequencies, i.e., before the onset of oscillatory behavior, do the simulation data deviate from the analytical approximation.
@@
Interestingly, as we approach the threshold, $a\tau/(\gamma b^{2})=1$, from above or below ($a=6k_{B}T$ and $a=25k_{B}T$) we see a significant deviation from the analytical prediction at small $\omega$, indicating that the simple linear approximation is no longer valid but a crossover behavior becomes relevant. However, the locations of the maxima and minima remain unaffected, from which we interpret that the signature of time delay persists even outside the linear regime.